OpenStudy (anonymous):

how do i integrate Sin(x)Cos(x)^2

OpenStudy (anonymous):

for this problem you need to use u substitution

OpenStudy (anonymous):

i thought so

OpenStudy (anonymous):

so should i set u=cosx^2 and du=-2sinx

OpenStudy (anonymous):

let \[u=\cos (x)\] so: \[du=-\sin(x) dx\]

OpenStudy (anonymous):

but its cos(x)^2

OpenStudy (anonymous):

\[\int\limits_{}^{} \sin(x) \cos^{2}(x) dx=-\int\limits_{}^{} u ^{2} du=-u^3/3=-(1/3)\cos^3(x)\]

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem